OpenStudy (anonymous):

For which operations is the set {–1, 0, 1} closed? Choose all answers that are correct. A. addition B. division C. multiplication D. subtraction

OpenStudy (anonymous):

I got Addition and subtraction for this question!!!! A. D.

OpenStudy (zzr0ck3r):

1+1 = 2 is not in the set -1-1 = -2 is not in the set

OpenStudy (zzr0ck3r):

1/0 is not in the set

OpenStudy (anonymous):

Thats correct yes u r right

OpenStudy (anonymous):

but by going up you need to use addition...

OpenStudy (zzr0ck3r):

-1*-1 = 1 -1*0 = 0 -1*1=-1 1*1=1 1*0=0 1*-1=-1 0*0=0 0*1=0 0*-1=0

OpenStudy (zzr0ck3r):

I don't understand what you mean. "by going up you need addition" please explain

OpenStudy (anonymous):

NVM could you go on with your explanation please

OpenStudy (zzr0ck3r):

well I showed you that it is not closed with respect to addition, and subtraction, and division. I did this by giving you counter examples Then I showed every possible case for multiplication, and in all cases we are still in the set.

OpenStudy (anonymous):

its multiple questions My friend

OpenStudy (anonymous):

jus so you know

OpenStudy (zzr0ck3r):

I don't know what you mean....

OpenStudy (anonymous):

so it definitely cant be Multipiclation cause its opened am i right??

OpenStudy (zzr0ck3r):

Is the set closed with respect to addition no is the set closed with respect to subtraction no is the set closed with respect to division no is the set closed with respect to multiplication yes

OpenStudy (anonymous):

Im sorry to say this but you have to choose more than one question

OpenStudy (zzr0ck3r):

A set \(S\) is closed with respect to an operation \(+\) provided that for all \(a,b\in S\) we have that \(a+b \in S\)

OpenStudy (zzr0ck3r):

What are you talking about?

OpenStudy (anonymous):

there is more than one answer...

OpenStudy (zzr0ck3r):

you are right. there are 4 no no no yes

OpenStudy (anonymous):

so there are three closed right or wrong just answer yes or no!

OpenStudy (anonymous):

what part of yes or no do you not understand

OpenStudy (zzr0ck3r):

wow dude listen I am helping you. The way you ask a question is not a yes or no answer.

OpenStudy (zzr0ck3r):

are you short or tall?

OpenStudy (zzr0ck3r):

yes or no?

OpenStudy (anonymous):

are there 3 closed yes or no

OpenStudy (zzr0ck3r):

there is only one set to be closed so when you say are they closed IT MAKES NO SENSE

OpenStudy (zzr0ck3r):

again that does not make sense.

OpenStudy (zzr0ck3r):

is the set closed with respect to 3 of the operations

OpenStudy (zzr0ck3r):

that is what you are asking

OpenStudy (zzr0ck3r):

the answer is no

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

Im sorry for being mean I'm just really tired

OpenStudy (zzr0ck3r):

I understand

OpenStudy (zzr0ck3r):

just listen for one second

OpenStudy (anonymous):

k

OpenStudy (zzr0ck3r):

if you take a set, and put an operation on it, then we call it closed if for any two elements in the set, if we apply the operation to it, the number we get in the end is also in the set

OpenStudy (zzr0ck3r):

so with addition 1 is in the set 1 is in the set but 1+1=2 is not in the set. So it is not closed

OpenStudy (anonymous):

so whats the answer

OpenStudy (zzr0ck3r):

but we do not call this open.

OpenStudy (zzr0ck3r):

it says choose the answers that are correct. the only one that is correct is D multiplication

OpenStudy (zzr0ck3r):

err C multiplication

OpenStudy (anonymous):

thats Impossible.... well I beleive u

OpenStudy (zzr0ck3r):

why?

OpenStudy (anonymous):

Can you help with more

OpenStudy (zzr0ck3r):

just first tell me why its impossible. it is obvious that you don't understand something fundamental, and going on is not going to help

OpenStudy (zzr0ck3r):

not trying to sound rude

OpenStudy (zzr0ck3r):

I don't understand the stuff I am working on either...that's how we learn

OpenStudy (anonymous):

well I took this assessment before I have 3 attempts the similar question to this was Multiple answers and Im pretty sure it wont change.

OpenStudy (anonymous):

It also tells you that Choose *ALL* anwsers hat are correct

OpenStudy (anonymous):

@zzr0ck3r were are u?!?

OpenStudy (zzr0ck3r):

why do you assume more than one is correct?

OpenStudy (anonymous):

I saw u answering billster's wuestion

OpenStudy (anonymous):

question

OpenStudy (zzr0ck3r):

again, if you understand what closed w.r.t. an operation is, then you would know I am right.

OpenStudy (anonymous):

It says all closed not open and Multiplication is open

OpenStudy (zzr0ck3r):

1+1=2 is NOT in the set and 1 is -1 - 1 = -2 is NOT in the set and 1,-1 are 1/0 is NOT in the set, and 1, 0 are then they are not closed

OpenStudy (zzr0ck3r):

there is no open

OpenStudy (zzr0ck3r):

either closed or not closed

OpenStudy (anonymous):

Just say all the closed one's so I understand please!!!

OpenStudy (zzr0ck3r):

D

OpenStudy (zzr0ck3r):

for the 1000th time

OpenStudy (zzr0ck3r):

ask anyone else here.

OpenStudy (anonymous):

OKAY yeesh!

OpenStudy (anonymous):

Which equations show that the set of whole numbers is not closed under subtraction? Choose all answers that are correct. A. 1 – (–2) = 3 B. 1 – 2 = –1 C. 2 – 0 = 2 D. 2 – 4 = –2

OpenStudy (zzr0ck3r):

I see a bunch of little boxs

OpenStudy (zzr0ck3r):

err not D on that last question C multiplication

OpenStudy (zzr0ck3r):

I keep thinking its D multiplication, but its C multiplication

OpenStudy (anonymous):

Which equations show that the set of whole numbers is not closed under subtraction? Choose all answers that are correct. A. 1 – (–2) = 3 B. 1 – 2 = –1 C. 2 – 0 = 2 D. 2 – 4 = –2

OpenStudy (anonymous):

refresh the page u should able to see it then!

OpenStudy (zzr0ck3r):

yep

OpenStudy (anonymous):

can u see now!

OpenStudy (zzr0ck3r):

yes

OpenStudy (zzr0ck3r):

ok what does it mean for a set to be closed?

OpenStudy (anonymous):

could you help please!

OpenStudy (zzr0ck3r):

I am. Answer my question

OpenStudy (zzr0ck3r):

I have explained this many times...

OpenStudy (anonymous):

If you can produce a number outside the set using the operation then it is the set is open under that operation. For example, integers are closed under addition because you can add any 2 integers together and ALWAYS get another integer.

OpenStudy (zzr0ck3r):

There is one more condition we must worry about. That is that the operation was done on two elements in our set to begin with.

OpenStudy (zzr0ck3r):

and its not called open, its called not closed.

OpenStudy (zzr0ck3r):

So @CavebusterWatcher! why is the first one bad? remember we must be applying the operation to elements in our set to get any meaningful result.

OpenStudy (anonymous):

guyz could you help me I will tell you if u r correct or not!

OpenStudy (zzr0ck3r):

I'll wait one more min for a reply that shows you gave thought and not trying to get the answer, but then I am gone.

OpenStudy (zzr0ck3r):

I know I am correct. We are here to help you learn, this is not helping if we give you answers, you put it in the computer, then you tell us we are right. All you learned to do is tell me something I already know.

OpenStudy (zzr0ck3r):

good luck. I would suggest trying to learn the material, the questions you are asking are very very easy.

OpenStudy (zzr0ck3r):

good night

OpenStudy (anonymous):

I have 5 question is this test its not much but you dont have to make this more complicated for me!!!!!!!!

OpenStudy (anonymous):

Wow what great help !!

OpenStudy (anonymous):

{NOT}

OpenStudy (anonymous):

Ok, I'll try to explain it a bit differently. Lets say I have a set of numbers {1,2,3,4,5,6,7} I also have operations that I can use on any to numbers to produce a result. For example one operation could be multiplication. Multiplication takes two numbers as an input and outputs another number. If a set is close under an operation, then applying an operation to any members of the set ALWAYS produces something that is already in the set. If it produces something not in the original set, then the set cannot be closed under that operation.

OpenStudy (anonymous):

My set is not closed under multiplication. Can you see how I can apply multiplication to 1,2,3,4,5,6,7 to produce a number that is not 1,2,34,5,6,7?

OpenStudy (zzr0ck3r):

We are not here to take your test. Read the rules. And later think about how someone you don't know spent about an hour with you trying to get you to understand something. Then think about how you gave it no thought at all and just wanted me to give you the answers. Then think about how rude you are to people like myself and @uhhhhhh then think about the grade you are getting in the class. Then ask yourself if there is a correlation between all these things. Once you have done this, and are ready to give it the effort it is going to take, you may tag me and I will gladly help you. Until then, good luck.